Grade 12Mechanics

A cubicle block of wood of edge 0.2m weighing 6.4kgf floats on water. Find the depth of immersion. Given, density of water 1000kg per cubic metre. Show working out please.

Myra Aseah
6 Years agoGrade 12

2 Answers

Arun
6 Years ago
s = 30 cm = 0.3 m
m = 21.6 kg
# Solution-
Let h be depth of immersion and Vi bs volume of immersed part.
Here, liquid displaced is equal to weight of wooden block.
mg = Vi.ρ0.g
m = s×s×h.ρ0
6.4 = 0.2^2 × 1000 × h
h = 0.16 m
h = 16 cm
 
Vikas TU
6 Years ago
Dear student 
30 × 30 × 10^-4 × h × g = 21.6
Let h = the depth of immersion then
H= 24
Good Luck 
Cheers